# 3D Gaussian Ray Tracing: Fast Tracing of Particle Scenes

### v2

@@ 7.1. Differences Between Ray Tracing and Rasterization
Although our ray tracing render achieves realtime performance, rasterization remains faster than ray tracing for the basic case of rendering primary rays from pinhole cameras (see  Table 2).
@@ Appendix A Implementation and Training Details
We wrap the NVIDIA OptiX tracer as a Pytorch extension and train our representation using Adam optimizer for $30000$ iterations. We set the learning rates for rotations, scales, and albedo to $0.001$, $0.005$, and $0.0025$, respectively. The learning rate for the remaining spherical harmonics coefficients is $20$ times smaller than that for albedo. Finally, we set the density learning rate to either $0.05$ for high-quality settings or $0.09$ for fast-inference settings.
After initial $500$ iterations, we start the densification and pruning process, which we perform until $15000$ iterations are reached. To densify the particles, we accumulate 3D positional gradients, scaled by half the distance of each particle to the camera, to prevent under-densification in distant regions. In line with 3DGS (Kerbl et al., 2023), we split the particles if their maximum scale is above $1\%$ of the scene extent and clone them otherwise. Pruning directly removes particles whose opacity is below $0.01$. Additionally, we employ a simple heuristic to cap the maximum number of particles to $3000000$. We denote this pruning strategy as visibility pruning. Specifically, if the densification step results in more particles, we reduce their number to $2700000$ by pruning particles with minimum accumulated weight contribution on the training views. Moreover, while densification and pruning are in effect and similar to 3DGS, we reset the particle densities to $0.01$ every $3000$ iterations. During training, we perform early stopping to terminate the tracing of rays whose accumulated transmittance falls below $0.001$. During inference, we increase this threshold to $0.03$ for improved efficiency. We begin by solely training albedo colors and progressively increase the spherical harmonics’ degree every $1000$ iterations, up to a maximum of $3$. We update the BVH every iteration and reconstruct it after each pruning and densification.

### v3

@@ 7.1. Differences Between Ray Tracing and Rasterization
For forward rendering, our approach achieves real-time performance, and is only about 2×2$\times$2 × slower than 3DGS’s tiled rasterization in the basic case of rendering primary rays from pinhole cameras (see  Table 2). For differentiable rendering to fit scenes, our ray tracing approach is 2×2$\times$2 ×-5×5$\times$5 × slower than rasterization, mainly due to the need to rebuild the BVH (Section  4.4), and additional arithmetic needed to evaluate the backward pass as terms are no longer shared between pixels. As an example, for the Tanks & Temples’s Truck scene, ray tracing is $\times3$ slower per iteration of optimization, with a mean iteration time of 100ms (30ms spent on the BVH construction, 15ms on the forward pass and 30ms on the backward pass), while the rasterization requires 30ms per-iteration (7ms on the forward pass and 10ms on the backward pass). Furthermore, our approach enables training with incoherent rays as discussed in Section 4.4, but in that case ray incoherency further increases the cost of ray tracing, leading to 5×5$\times$5 × slower optimization.
@@ 7.2. Limitations and Future Work
Acknowledgements.
We are grateful to Hassan Abu Alhaija, Ronnie Sharif, Beau Perschall and Lars Fabiunke for assistance with assets, to Greg Muthler, Magnus Andersson, Maksim Eisenstein, Tanki Zhang, Dietger van Antwerpen and John Burgess performance feedback, to Thomas Müller, Merlin Nimier-David, and Carsten Kolve for inspiration, to Ziyu Chen, Clement Fuji-Tsang, Masha Shugrina, and George Kopanas for experiment assistance, and to Ramana Kiran and Shailesh Mishra for typo fixes. The manta ray image is courtesy of abby-design.
@@ Appendix A Implementation and Training Details
We wrap the NVIDIA OptiX tracer as a Pytorch extension and train our representation using Adam optimizer for $30000$ iterations. We set the learning rates for rotations, scales, and zeroth-order spherical harmonics to $0.001$, $0.005$, and $0.0025$, respectively. The learning rate for the remaining spherical harmonics coefficients is $20$ times smaller than for the zeroth-order coefficient. Finally, we set the density learning rate to either $0.05$ for high-quality settings or $0.09$ for fast-inference settings.
After initial $500$ iterations, we start the densification and pruning process, which we perform until $15000$ iterations are reached. To densify the particles, we accumulate 3D positional gradients, scaled by half the distance of each particle to the camera, to prevent under-densification in distant regions. In line with 3DGS (Kerbl et al., 2023), we split the particles if their maximum scale is above $1\%$ of the scene extent and clone them otherwise. Pruning directly removes particles whose opacity is below $0.01$. Additionally, we employ a simple heuristic to cap the maximum number of particles to $3000000$. We denote this pruning strategy as visibility pruning. Specifically, if the densification step results in more particles, we reduce their number to $2700000$ by pruning particles with minimum accumulated weight contribution on the training views. Moreover, while densification and pruning are in effect and similar to 3DGS, we reset the particle densities to $0.01$ every $3000$ iterations. During training, we perform early stopping to terminate the tracing of rays whose accumulated transmittance falls below $0.001$. During inference, we increase this threshold to $0.03$ for improved efficiency. We begin by solely training the constant spherical harmonic and progressively increase the spherical harmonics’ degree every $1000$ iterations, up to a maximum of $3$. We update the BVH every iteration and reconstruct it after each pruning and densification.